# CricBench: A Multilingual Benchmark for Evaluating LLMs in Cricket Analytics

### v1

@@ Abstract
Cricket is the second most popular sport globally, commanding a massive following of over 2.5 billion fans globally. Enthusiasts and analysts frequently seek advanced statistical insights—such as long-term historical performance trends or complex player comparisons—that are often unavailable through standard web searches. While Large Language Models (LLMs) have advanced significantly in Text-to-SQL tasks, their capability to handle the domain-specific nuances, complex schema variations, and multilingual requirements inherent to sports analytics remains under-explored. To investigate this potential capability gap, we present CricBench, a comprehensive benchmark suite for evaluating LLMs on specialized cricket data. To curate a “Gold Standard” dataset, we collaborate with domain experts in cricket and SQL to manually author complex queries, ensuring logical correctness. Recognizing linguistic diversity, we construct the benchmark in both English and Hindi, establishing a framework that is open for further extension to other regional languages. We evaluate six state-of-the-art models—including GPT-4o, Claude 3.7 Sonnet, and open-source models using a strict evaluation protocol. Our results reveal that high performance on general benchmarks does not guarantee success in specialized domains. While the open-weights reasoning model DeepSeek R1 achieves state-of-the-art performance (50.6%), surpassing proprietary giants like Claude 3.7 Sonnet (47.7%) and GPT-4o (33.7%), it still exhibits a significant accuracy drop when moving from general benchmarks (BIRD) to CricBench. Furthermore, we observe that code-mixed Hindi queries frequently yield parity or higher accuracy compared to English, challenging the assumption that English is the optimal prompt language for specialized SQL tasks.
@@
Vaibhav Devraj, Dhruv Kumar, Jagat Sesh Challa Birla Institute of Technology and Science (BITS), Pilani f20211107@pilani.bits-pilani.ac.in {dhruv.kumar, jagatsesh}@pilani.bits-pilani.ac.in
@@ 1 Introduction
Large Language Models (LLMs) have achieved significant milestones in translating natural language into structured query language (Text-to-SQL), fundamentally altering how non-technical users interact with databases Chen et al. (2021); Achiam et al. (2023). Pioneering benchmarks such as WikiSQL Zhong et al. (2017) and Spider Yu et al. (2018) have driven rapid improvements in schema linking and syntactic correctness Wang et al. (2020). More recently, the BIRD benchmark Li et al. (2023) has pushed the frontier by integrating external knowledge reasoning. However, while general-purpose models excel at standard business queries, the extent to which these capabilities transfer to highly specialized domains remains an open research question Suhr et al. (2020).
@@
Current Text-to-SQL benchmarks are predominantly English-centric Yu et al. (2018); Li et al. (2023). This limitation is particularly critical for sports like cricket, where the fanbase is linguistically diverse. In regions like India, users often interact in native languages or code-mixed dialects. A model that generates accurate SQL for an English query but fails to process the same intent in Hindi is of limited utility in real-world deployment He et al. (2019).
@@
  • CricBench Dataset: We release the first expert-curated, domain-specific Text-to-SQL benchmark for cricket analytics, consisting of normalized schemas and high-complexity queries. [^1] The dataset and code are publicly available at https://github.com/Vaibhav2823/ipl
@@ 2.2 Domain-Specific Query Generation
Research has increasingly demonstrated that generalist models often struggle with the nuances of specialized domains like healthcare and finance, where "correct" SQL generation depends heavily on understanding domain logic rather than just schema linking Suhr et al. (2020). Previous works have attempted to mitigate this via Retrieval-Augmented Generation (RAG) Lewis et al. (2020) and few-shot prompting Brown et al. (2020). However, standard retrieval often misses strict logical constraints (e.g., specific calculation formulas). Our work extends this by implementing a deterministic routing layer. Instead of retrieving unstructured context, our system injects precise domain rules (e.g., franchise evolution, match phases) to fill the parametric knowledge gap, ensuring the model can handle details that were not present in its initial training data. Additionally, we employ a rigorous manual authoring process by domain experts Wei et al. (2022) to ensure high data fidelity.
@@ 2.3 Multilingual Evaluation
While English-centric Text-to-SQL research is mature, multilingual benchmarks remain scarce for Indic languages. Existing resources often focus on high-resource languages or rely on machine-translated versions of general benchmarks He et al. (2019). Our work addresses this gap by introducing a verified Hindi test set. We specifically address the challenge of code-mixing, where technical terms are retained in English (e.g., "Strike Rate") within Hindi sentence structures, mirroring authentic user behavior in the Indian subcontinent.
@@ 7 Data Availability
To facilitate future research in domain-specific Text-to-SQL, we make all artifacts from this study publicly available.
@@
The release package includes:

### v2

@@ Abstract
Cricket is the second most popular sport globally, commanding a massive following of over 2.5 billion fans globally. Enthusiasts and analysts frequently seek advanced statistical insights such as long-term historical performance trends or complex player comparisons that are often unavailable through standard web searches. While Large Language Models (LLMs) have advanced significantly in Text-to-SQL tasks, their capability to handle the domain-specific nuances, complex schema variations, and multilingual requirements inherent to sports analytics remains under-explored. To investigate this potential capability gap, we present CricBench, a comprehensive benchmark suite for evaluating LLMs on specialized cricket data. To curate a "Gold Standard" dataset, we collaborate with domain experts in cricket and SQL to manually author complex queries, ensuring logical correctness. Recognizing linguistic diversity, we construct the benchmark in both English and Hindi, establishing a framework that is open for further extension to other regional languages. We evaluate six state-of-the-art models—including GPT-4o, Claude 3.7 Sonnet, and open-source models using a strict evaluation protocol. Our results reveal that high performance on general benchmarks does not guarantee success in specialized domains. While the open-weights reasoning model DeepSeek R1 achieves state-of-the-art performance (50.6%), surpassing proprietary giants like Claude 3.7 Sonnet (47.7%) and GPT-4o (33.7%), it still exhibits a significant accuracy drop when moving from general benchmarks (BIRD) to CricBench. Furthermore, we observe that code-mixed Hindi queries frequently yield parity or higher accuracy compared to English, challenging the assumption that English is the optimal prompt language for specialized SQL tasks.
@@
Vaibhav Devraj, Dhruv Kumar, Jagat Sesh Challa, Parth Agarwal, Navya Kommuri, Trizal Garg, Prisha Singhal, and Dhruv Shah Birla Institute of Technology and Science (BITS), Pilani {f20211107, dhruv.kumar, jagatsesh, f20240726, f20240594, f20240338, f20240515, f20240529}@pilani.bits-pilani.ac.in
@@ 1 Introduction
Large Language Models (LLMs) have achieved significant milestones in translating natural language into structured query language (Text-to-SQL), fundamentally altering how non-technical users interact with databases Chen et al. (2021); Achiam et al. (2023). Pioneering benchmarks such as WikiSQL Zhong et al. (2017) and Spider Yu et al. (2018) have driven rapid improvements in schema linking and syntactic correctness Wang et al. (2020). More recently, the BIRD benchmark Li et al. (2023) has pushed the frontier by integrating external knowledge reasoning. However, while general-purpose models excel at standard business queries, the extent to which these capabilities transfer to highly specialized domains remains an open research question Suhr and others (2020).
@@
Current Text-to-SQL benchmarks are predominantly English-centric Yu et al. (2018); Li et al. (2023). This limitation is particularly critical for sports like cricket, where the fanbase is linguistically diverse. In regions like India, users often interact in native languages or code-mixed dialects. A model that generates accurate SQL for an English query but fails to process the same intent in Hindi is of limited utility in real-world deployment He and others (2019).
@@
  • CricBench Dataset: We release the first expert-curated, domain-specific Text-to-SQL benchmark for cricket analytics, consisting of normalized schemas and high-complexity queries. [^1] The dataset and code will be made publicly available upon acceptance.
@@ 2.2 Domain-Specific Query Generation
Research has increasingly demonstrated that generalist models often struggle with the nuances of specialized domains like healthcare and finance, where "correct" SQL generation depends heavily on understanding domain logic rather than just schema linking Suhr and others (2020). Previous works have attempted to mitigate this via Retrieval-Augmented Generation (RAG) Lewis et al. (2020) and few-shot prompting Brown et al. (2020). However, standard retrieval often misses strict logical constraints (e.g., specific calculation formulas). Our work extends this by implementing a deterministic routing layer. Instead of retrieving unstructured context, our system injects precise domain rules (e.g., franchise evolution, match phases) to fill the parametric knowledge gap, ensuring the model can handle details that were not present in its initial training data. Additionally, we employ a rigorous manual authoring process by domain experts Wei et al. (2022) to ensure high data fidelity.
@@ 2.3 Multilingual Evaluation
While English-centric Text-to-SQL research is mature, multilingual benchmarks remain scarce for Indic languages. Existing resources often focus on high-resource languages or rely on machine-translated versions of general benchmarks He and others (2019). Our work addresses this gap by introducing a verified Hindi test set. We specifically address the challenge of code-mixing, where technical terms are retained in English (e.g., "Strike Rate") within Hindi sentence structures, mirroring authentic user behavior in the Indian subcontinent.
@@ 7 Data Availability
To facilitate future research in domain-specific Text-to-SQL, we will make all artifacts from this study publicly available upon acceptance.
@@
The release package will include: